the sum of two numbers is 12. three times the larger minus 4 is equal to 5 times the smaller. what are the numbers?
Answer by checkley71(8403)

x+y=12 or x=12-y
3y-4=5x
3y-4=5(12-y)
3y-4=60-5y
3y+5y=60+4
8y=64
y=64/8
y=8 answer.
x=12-8
x=4 answer.
proof:
3*8-4=5*4
24-4=20
20=20